OpenStudy (anonymous):

If someone can help me with this that would be greatly appreciated, but if no one can that's okay. Related Rate Problem: A hemispherical tank with a radius of 10 m is filled from an inflow pipe at a rate of 3 m^3/min. (Hint: The volume of a cap of thickness h sliced from a sphere of radius r is ∏h^2(3r-h)/3) a. how fast is the water level rising when the water level is 5 m from the bottom of the tank? b. what is the rate of change of the surface area of the water when the water is 5 m deep?

OpenStudy (anonymous):

When depth is h, volume in cap = V = πh²(30−h)/3 = 10πh² − πh³/3 The rate at which the water level is rising = dh/dt. To find it we differentiate above with respect to t. Note for this purpose both V & h are regarded as functions of t and so we use the chain rule “ df/dt = f'.dh/dt “ to get the t-derivative of any function f of h. dV/dt = (20πh − πh²).dh/dt The question gives dV/dt=3 when h=5 so we can use this equation to get dh/dt. 3 = (100π−25π).dh/dt → dh/dt = 1/(25π) = 0.0127 m/min (b) If we can find a formula connecting surface area S and h then we can differentiate again. This will gives us a link between dS/dt and dh/dt. Since we now know dh/dt, dS/dt will follow. Draw a diagram of the central vertical cross-section and apply Pythagoras to find the radius R of the surface disc. 10² = R² + (10−h)² → R² = 20h−h² Surface area of water, S = πR² = π(20h−h²) Differentiate wrt t : dS/dt = π(20−2h).dh/dt = π*10*{1/(25π)} = 0.4 m²/min ( I used the π version of dh/dt because it was obvious the π would cancel )

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

yeah! i just need to read over it a couple of times lol. thank you so much for your help!